exklusiveyanna:

\frac{f+3}{12}=\frac{7}{2} cross multiplying

Vocaloid:

\[\frac{ f+3 }{ 12 }=\frac{ 7 }{ 2 }\] when we cross multiply, we multiply in a cross direction like this: 

Vocaloid:

to finish solving this, distribute the 2 across f+3, then solve for f